The price in a market is dominated by two firms is affected by the quantities supplied by both firms, Q1 and Q2: P = 150 - (Q1 + Q2). The marginal cost for the two firms is identical and constant and equal to 20.

a. Derive the equations for total revenue for the two firms.
b. Compute the profit-maximizing levels of output and prices for the firms.
c. Compute the profit-maximizing level of output and price for the industry if the duopolists merged and formed a monopoly.
d. Compare and contrast the results from b. and c.
Answer #1

The prices in a market is two firms ii affected by both | Q2; p= 150-(Qc+Q2] les in a market is dominated by o e affected byVeu Total revenue for firm 2 TR2 = 150 Q2 - Q1 Q2 - Q2 Marginal reve me for füm 2. MR2 = a (TR2) = 150-Q1-2Q2 Now, Marginal clevel of output for at manimising : 43.3 and Hnce profit manimising lead of both tims Q = 43.3 and Q2²4 Prices for both firmsje 65 units. Also, in duopoly case, Price charged by both birms were much lown (ie $63.4) than the price charged by the .. mo